OpenStudy (cookiemonster18):

Tell whether the lines for each pair of equations are parallel, perpendicular, or neither. y = -1/4x + 10 –2x + 8y = 6 (1 point) A.parallel B.perpendicular C.neither I think the answer is C (neither) am I correct?

OpenStudy (cookiemonster18):

@Vocaloid can you help me :(

OpenStudy (anonymous):

For both lines, can you tell me what the gradient is please?

OpenStudy (danjs):

parallel lines will have th esame rate of change, or slope, or direction perpendicular lines will have a slope that is negative and inverted

OpenStudy (danjs):

slope m perpendicular - slope -1/m

OpenStudy (danjs):

change the second one to look like the first y=m*x + b y = 2/8 x + 6/8 ----> y = 1/4 x + 3/4 same slope 1/4 as the first line